OpenStudy (anonymous):

A postal clerk sold 75 stamps for $4.80. Some were 5-cent stamps and others were 8-cent stamps. How many of each type of stamp did he sell?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

let x = 5 cent stamps and y = 8 cent stamps, youll get these equations: x + y = 75 which is the number of stamps 5x + 8y = 480 which is the value of the stamps.

OpenStudy (anonymous):

if you still cant solve them, x = 75 - y, sub in 2nd eq, 5(75-y) + 8y = 480 375 + 3y = 480 3y = 105 y= 35 then sub back into 1st eq, x + 35 = 75 x = 40